EUROPEAN MATHEMATICAL SOCIETY

Report on the Meeting of the Executive Committee Besançon (France) October 20-21, 1995

Eva Bayer Fluckiger, Jean-Pierre Bourguignon, Alberto Conte, Aatos Lahtinen, László Márki, Peter Michor, Andrzej Pelczar, V.A. Solonnikov and David Wallace were present. Isabel Labouriau was excused.

Pierre Bérard (Chairperson of the Committee on Developing Countries), Mireille Chaleyat-Maurel (Public Relations Officer), Luc Lemaire (Liaison Officer with the European Union), Bernard Prum (Chairperson of the Round Tables Committee of ECM2), Jean-Pierre Puel (President of the Société de Mathématiques Appliquées et Industrielles (SMAI)), Stewart Robertson (Chairperson of the Publications Committee), Tuulikki Mäkelainen (Secretary of the European Mathematical Society) were invited.

The Executive Committee thanks Eva Bayer for the perfect arrangements and hospitality during the meeting.

SCIENTIFIC ACTIVITIES

European Congress of Mathematics

ECM2 July 22-26, 1996 Budapest (Hungary) See the Second Announcement in the Newsletter.

Diderot Mathematical Forum

The first conference of the series will take place in the autumn of 1996 with the theme “Mathematics and Finance”. It will be organized by P.Embrechts (Zurich) with the following towns and respective local organizers: London (M.Davis), Moscow (A.Shiryaev), Zurich (H.Bühlmann).

The conference could have a day in common in Zurich with the International Statistical Institute which had the same project of a conference on mathematical finance.

Journal of the European Mathematical Society (JEMS)

The special Committee has started to negotiate with a potential Editor-in-Chief. A small group (maybe four) of editors and a larger editorial board are being considered. There should be no restrictions as to the assistant editors but they should be drawn from a balanced list, and the Editor in Chief should have a decisive role in the choice of the assistant editors.

CENTRAL AND EASTERN COUNTRIES COMMITTEE

The difficulty of travelling for Eastern Europe mathematicians and the situation of libraries in Central and Eastern Europe (CEE) are the major concerns.

The Committee suggests that its next action in this direction should be targeted, limited in the time, and closer to the activity of EMS. The Executive Committee thinks that efforts in 1996 should be concentrated on activities around the Budapest Congress and its satellites (information on satellite conferences is on the servers).

As regards libraries, a letter has been sent to some identified institutes. The first output of this action will be presented later on.

COMMITTEE ON DEVELOPING COUNTRIES

The European Union and also national governments have to be persuaded to see the needs of mathematics in developing countries.

One problem is that of libraries. Photocopying services for mathematicians in several developing countries are on demand. Centres with online Zentralblatt could allow developing countries to use their service; but there are problems with personnel time, mailing costs and copyright.

IMU-CDE plans a directory of mathematics (departments etc.) in different developing countries.

The Committee should contact EMS member societies to find out about their programs concerning developing countries.

Education of teachers is crucial for developing countries, having in mind that it needs to be adapted to the local society. There are different classes of developing countries: ones with active mathematical research have also good mathematical education, others lack even school mathematics education.

During the African Mathematical Union (AMU) meeting in Ifrane, it was discussed that African countries wish to join European networks. South Africa joined AMU; they have launched a big change in education with interesting ideas. The North African group would like to have closer connections with EMS. A directory of Centres of Excellence in Mathematics and of Mathematicians in Third World Countries should be very useful: ACMS, ICTP, CIMPA, IMU CDE, and EMS have to cooperate.

The composition of the Committee on Developping Countries is the following : Chairperson : P. Bérard, members: K.D. Elworthy, M. Giaquinta, B. Goldsmith, J.P. Gossez, J.F. Jaulent, V.N. Maslennikova, A. Pelczar, J.F. Rodrigues and G. Schiffels.

RELATIONS WITH EUROPEAN INSTITUTIONS

Group on Relations with European Institutions

Information gathering and distribution have been successful, advice on making applications has also been well received. The number and the level of applications in mathematics has risen. Regular contacts have been established in Brussels, at different levels. The European Commission has a server ( http://www.cordis.lu). A link to this server should be made available on the EMS-server.

The changes in programmes are the following :

⋄ “Training and Mobility of Researchers” instead of “ Human Capital and Mobility”, which is discontinued.

⋄ “Large Scale Facilities” : most of the money goes to facilities, not to fellows, which is problematic for mathematics.

⋄ Networks: new rules have been installed to accommodate the new policy: institutional fellowships abolished, and the money moved to networks, each network to have at least 5 laboratories from 3 different countries with a budget of at least 1 MEcu. There is need for flexible rules so that for example the mathematical panel will be allowed to attribute smaller sums per network. Another possibility would be to allow the panel to merge some applications of scientifically close networks.

⋄ Euroconferences and Summer Schools : The programme functions now, with a definite policy: about 100 persons attending conferences organized by European oriented societies. It is suggested to apply summer schools with money to Eastern European mathematicians as travel grants.

The EMS has to think about an influence on the new budget for Eastern countries and the national representatives/committees in Brussels.

Lobbying in Brussels is necessary. Cooperation with people in industry is needed. EMS should get other societies to be corporate members.

The Committee on the Relations with European Institutions is renamed the Group on Relations with European Institutions with the following composition: Chairperson: J.-P. Bourguignon, with members : E. Bayer, L. Lemaire, F. van Oystaeyen.

Preparations for the Fifth Framework Programme

An EMS-ESF-ESTA (European Science and Technology Assembly) task force met once with the following composition : A. Bensoussan, J.-P. Bourguignon, M. Espedal, J. Fenstad, K.H. Hoffmann and H.P.F. Swinnerton-Dyer. It worked fruitfully and produced a draft in order to make mathematics more present in the Programmes of the European Commission. It pointed out that the future development of mathematics can be severely handicaped by policy makers’ inadequate appreciation of three basic facts :

⋄ the rate at which new mathematics is produced continues unabated.

⋄ the impact that mathematics has on society is growing so fast and so steadily that the number of contact points between mathematics and the man in the street has gone up considerabily without his noticing it.

⋄ the need for keeping all components of this science in connection with one another has never been as strong as it is now.

Euresco The European Research Conference programme, conducted by the ESF, is a series of conferences, every

other year, on topics of the highest interest and importance. The aim is to hold finite series. Funds are basically from EU, imposing restrictions. Speakers are invited, young people are encouraged to apply. Information on the programme is to be announced in the Newsletter. The conferences should not be summer schools.

Next conferences are : European Research Conferences, Mathematics: ⋄ Mathematical Methods in Industrial Problems 1996 : Multi–scale Analysis in Image Processing: Mathematical and Numerical Models for the Simu-

lation of Turbulent and Reactive Flows, Alfio Quarteroni (Milano), February 1996, Mont St. Odile (near Strasbourg), France.

⋄ Algebra and Discrete Mathematics 1996 : Group Theory: Finite to Infinite (Laszlo Babai, Chicago), Castelvecchio, Italy, 13–18 July, 1996 ⋄ Number Theory and Arithmetical Geometry 1997 : (Gerhard Frey, Essen), autumn of 1997.

INFORMATION SERVICES

European Mathematical Information Service (EMIS) The EMS server (EMIS) is functioning at http://www.emis.de. Its launching had a great success :

about 600 people got connected on the first day.

Each individual member has to be given an opportunity to refuse to have their name and address in a public directory accessible through EMIS. A personal letter to individual members asking them to indicate if they refuse to have their name and address on the server shall be sent. The individual members are given a code to have access to the list for a limited period of time, in order to supply the individual members a possibility to check the members for the elections of delegates of individual members.

WORLD MATHEMATICAL YEAR 2000 (WMY 2000)

V. L. Hansen is appointed as the chairperson of the WMY 2000 Committee of EMS. The EMS Committee is composed as follows: Mireille Chaleyat-Maurel (France), Alberto Conte (Italy),

Stephen Gelbart (Israel), Vagn Lundsgaard Hansen (Denmark), Michael Hazewinkel (Netherlands), Osmo Pekonen (Finland), Jose Francisco Rodrigues (Portugal).

The member Societies should be informed about the plans for the World Mathematical Year 2000.

RELATIONS WITH MATHEMATICAL INSTITUTIONS

Tbilissi International Centre (Georgia) The three representatives of EMS appointed on the board of the Centre are : J. C. Hausmann (Switzer-

land), G. Roach (Scotland) and S. Solonnikov (Russia). SOCIETY’S LIFE

EMS Council

Elections to the Council The elections of individual member’s delegates will take place as follows : ⋄ December, 1995 : nominations are called for in the Newsletter. ⋄ End of February, 1996 : nomination forms must arrive at the Society’s Office in Helsinki. ⋄ March, 1996 : postal elections by the individual members.

Council Meeting The Council Meeting of the EMS will take place on July 20-21, 1996 in Budapest. The delegates should

arrive on Friday, July 19, 1996 and a (modest) reception should be organized so that the council delegates have the opportunity to meet members of the Executive committee. The Council Meeting is to begin at 9 a.m. on Saturday and the venue is the Lecture Room of the Hungarian Academy of Sciences.

Charta for electronic publishing will be added to the agenda, if ready.

Relationship with member societies EMS Vice-President D. Wallace will send a letter to corporate members in order to build a database

and to get a better knowledge of both present ones as well as candidates. Societies are asked to mention that they belong to EMS in their communications with the following points :

⋄ application for individual membership should made easy and be publicized. ⋄ information on ECM2 and WMY 2000. ⋄ the address of each Society’s server is to be obtained, if there is one. ⋄ individual members will be approached on whether they agree to have their name and address on the

server EMIS.